# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 1.1. Problem formulation
In the same way that Corollary 1.1 was proved, we can extend Corollary 1.2 to establish the finite sample memorization property for labels in $\mathbb{R}^{m}$, in this case, using a neural network of width $2m$, and a depth of $2N+4M$ layers.
@@
We now analyze the property of universal approximation. Our second main result reads as follows:
@@
$\|\phi^{\mathcal{L}}(\mathcal{W}^{\mathcal{L}},\mathcal{B}^{\mathcal{L}},\cdot)-f(\cdot)\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon.$ (1.6)
@@
$\mathcal{L}(\varepsilon)\leq C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{dp}\varepsilon^{-dp},$ (1.7)
@@
Theorem 1.2 ensures the existence of a neural network with a fixed width of $d+1$ neurons and sufficient depth approximating any function in $L^{p}(\Omega;\mathbb{R}_{+})$. As illustrated in the strategy of proof below, the neural network is constructed using geometrical arguments.
@@
Step 1. Let $f\in L^{p}(\Omega;\mathbb{R}_{+})$ be a given function and $\varepsilon>0$. Denote by $\mathcal{C}$ the smallest hyperrectangle containing $\Omega$, oriented according to the axes of the canonical basis of $\mathbb{R}^{d}$. We extend $f$ by zero into $\mathcal{C}$. Then, we construct a particular simple function $f_{h}$ that approximates $f$. For its construction, we consider a family $\mathcal{H}_{h}$ of hyperrectangles of size $h>0$ such that $\mathcal{H}_{h}\cup G_{h}^{\delta}=\mathcal{C}$, where $G_{h}^{\delta}$ is a grid with thickness $\delta<h$ that satisfies $m_{d}(G_{h}^{\delta})\rightarrow 0$ as $\delta\rightarrow 0$. Then, $f_{h}$ is defined as the average value of the function $f$ on each hyperrectangle of the family $\mathcal{H}_{h}$. With this simple function, we can guarantee that there exists $h_{1}>0$ such that for every $h<h_{1}$, we have $\|f-f_{h}\|_{L^{p}(\mathcal{C};\mathbb{R}{+})}<\varepsilon/2$. We denote by $N_{h}$ the number of hyperrectangles in $\mathcal{H}_{h}$.
Step 2. Then, we construct a neural network $\phi^{\mathcal{L}}$ with a width $d+1$ and depth $\mathcal{L}$, ensuring the existence of $\delta_{1}>0$ such that for all $\delta<\delta_{1}$, we have $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(\mathcal{C};\mathbb{R}_{+})}<\varepsilon/2$. This is done in two steps:
Step 2.1. We consider $\mathcal{H}^{E}_{h}$ to be a subset of hyperrectangles in $\mathcal{H}_{h}$ that are closest to the edges of $\mathcal{C}$, with $N^{E}_{h}$ denoting the number of hyperrectangles in $\mathcal{H}^{E}_{h}$. For each hyperrectangle $H\in\mathcal{H}^{E}_{h}$, we construct a two-layer neural network with an input-output map $\phi_{1}^{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. This map, which has a width of $d+1$, drives $H$ to a single point in $\mathbb{R}^{d}$ while mapping the remaining hyperrectangles in $\mathcal{H}_{h}$ to distinct, non-overlapping hyperrectangles in different locations. Here, the chosen parameters ensure the injectivity of the neural network with respect to the hyperrectangles in $\mathcal{H}_{h}$.
@@
Step 2.2 Then, we apply Theorem 1.1 to construct an input-output map $\phi^{L}$ that drives the points $\{x_{i}\}_{i=1}^{N_{h}}$ to the $M_{h}$ values of $f_{h}$, ensuring that the composition $\phi^{\mathcal{L}}:=(\phi^{L}\circ\phi^{2N^{E}_{h}})$ equals $f_{h}$ on $\mathcal{H}_{h}$, and therefore $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(\mathcal{H}_{h};\mathbb{R}_{+})}=0$. Additionally, we estimate the error introduced by the neural network on $G_{h}^{\delta}$ and show that for a fixed $h<1$, the norm $\|\phi^{\mathcal{L}}\|_{L^{\infty}(G_{h}^{\delta};\mathbb{R}_{+})}$ is bounded. This estimation allows us to ensure that there exists a $\delta_{1}>0$ such that for all $\delta<\delta_{1}$, we have $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(G_{h}^{\delta};\mathbb{R}_{+})}<\epsilon/2$.
Step 3. As a consequence of the triangle inequality and by choosing $\delta<\min{h_{1},\delta_{1}}$, we obtain $\|f-\phi^{\mathcal{L}}\|_{L^{p}(\Omega;\mathbb{R}{+})}<\varepsilon$.
Step 4. Due to the explicit construction of the neural network $\phi^{\mathcal{L}}$, the depth $\mathcal{L}$ can be estimated in terms of $h$ and $\delta$. Additionally, assuming $W^{1,p}$ regularity of $f$, explicit estimates for $h_{1}$ and $\delta_{1}$ can be obtained in terms of $\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}$, $\varepsilon$, $p$ and $d$ (see [12] and its application to isotropic partitions), and therefore conclude (1.7).
@@
$\bullet$ We emphasize that the most challenging step of the proof is step 2, which involves the neural network approximation.
$\bullet$ As shown in Section 2, each neuron in the neural network represents a hyperplane. In step 2.1, $d+1$ neurons are necessary because this is the number of hyperplanes required to separate a single hyperrectangle in $\mathcal{H}^{E}_{h}$. For instance, in the 2-dimensional case, a hyperrectangle located on the left edge of $\mathcal{C}$ has only one adjacent hyperrectangle above, one below, and one to its right (there are no hyperrectangles outside $\mathcal{C}$). Thus, only $d+1=3$ hyperplanes are necessary to separate this hyperrectangle from the others, as opposed to the case where a hyperrectangle is in the interior of $\mathcal{C}$, where four hyperrectangles would surround it, and it would be necessary to use $42d=$ 4 hyperplanes. These $d+1$ hyperplanes, in particular, ensure the compression of the separated hyperrectangle into a point.
$\bullet$ The set $\mathcal{H}^{E}_{h}$ allows us to define a neural network with $2N^{E}_{h}$ layers, mapping the $N_{h}$ hyperrectangles of $\mathcal{H}_{h}$ to $N_{h}$ distinct points. Focusing on $\mathcal{H}^{E}_{h}$ is essential, as it reduces the number of required layers. Mapping each hyperrectangle of $\mathcal{H}_{h}$ individually would result in a neural network with significantly more layers, i.e., $N_{h}\gg N^{E}_{h}$.
$\bullet$ The neural network in Theorem 1.2, defined as $\phi^{\mathcal{L}}=\phi^{L}\circ\phi^{2N_{h}^{E}}$ requires a width of $d+1$. However, only the first $2N^{E}_{h}$ layers need to maintain this width, while the width of the $L$ remaining layers can be reduced to $2$ (as a consequence of Theorem 1.1).
$\bullet$ The neural network of width $2$ from Theorem 1.1 maps the compressed hyperrectangles to their corresponding labels and requires $L=2N_{h}+4M_{h}-1$ layers, where $M_{h}$ is the number of distinct values taken by the approximating simple functions $f_{h}$ (which act as labels). In particular, $M_{h}\leq N_{h}$, and $N_{h}$ can be estimated in terms of $h$ and $\delta$. Moreover, since $N^{E}_{h}\leq N_{h}$ and $M_{h}\leq N_{h}$, the depth of the neural network $\phi^{\mathcal{L}}$ is bounded by $\mathcal{L}=2N^{E}_{h}+2N_{h}+4M_{h}-1\leq 8N_{h}-1$. Finally, following Step 4, we can estimate the total depth.

### v2

@@ 1.2. Problem formulation
Following the same approach used in the proof of Corollary 1.2, we can extend Corollary 1.3 to establish the finite sample memorization property for labels in $\mathbb{R}^{m}$. In this case, the construction yields a neural network of width $2m$ and depth $2N+4M$.
1.3.2. Implications for Neural Network Training
The training of neural networks is typically formulated as the minimization of a regularized empirical risk functional. Given a dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\mathbb{R}^{m}$, the standard training objective reads
$\min_{(\mathcal{W}^{L},\mathcal{B}^{L})}\left\{\mathcal{J}_{\lambda}(\mathcal{W}^{L},\mathcal{B}^{L})=\lambda\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\phi(\mathcal{W}^{L},\mathcal{B}^{L},x_{i}),y_{i}\right)\right\},$ (1.9)
where the first term penalizes large weights with a parameter $\lambda>0$, and the second term measures the prediction error on the training dataset through the input-output map $\phi(\mathcal{W}^{L},\mathcal{B}^{L},\cdot)$ defined in (1.2). Here, $\operatorname{loss}(\cdot,\cdot)$ is a given continuous and nonnegative function such that $\operatorname{loss}(x,x)=0$. For instance, $\operatorname{loss}(x,y):=\|x-y\|^{2}_{p}$ with $p\in\{1,2\}$ is commonly used for regression tasks, while $\operatorname{loss}(x,y)=\log(1+e^{x})-yx$, with $y\in\{0,1\}$ and $x\in\mathbb{R}$, is used for binary classification (binary logistic loss). The regularization parameter $\lambda$ controls the trade-off between data fidelity and model complexity.
Note that for every $\lambda>0$, the functional $\mathcal{J}_{\lambda}$ is coercive. Moreover, since both $\operatorname{loss}$ and the activation functions $\bm{\sigma}_{j}$ are continuous, the Bolzano–Weierstrass theorem ensures the existence of optimal parameters for every $\lambda>0$.
Recall that Theorem 1.1 establishes the existence of $(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$ such that exact interpolation of arbitrary finite datasets is possible. In particular, this guarantees that
$\mathcal{J}_{0}(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})=0,$ (1.10)
and since $\mathcal{J}_{0}\geq 0$, we deduce that $(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$ is a minimizer of $\mathcal{J}_{0}$. Moreover, we have the following result:
Theorem 1.2.
Let $d,\,N,\,m\in\mathbb{N}$ with $d,\,N,\,m>1$, and let $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\mathbb{R}^{m}$. Let $(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})$ be a minimizer of $\mathcal{J}_{\lambda}$, and let $(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$ be a minimizer of $\mathcal{J}_{0}$ for the ReLU activation function. Then, we have
$\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\leq\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2},\quad\text{for all }\lambda>0.$ (1.11)
In particular, we have
$\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\phi(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda},x_{i}),y_{i}\right)\rightarrow 0\quad\text{as }\lambda\rightarrow 0,$ (1.12)
and
$\left|\!\left|\!\left|(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})\right|\!\right|\!\right|_{2}\leq\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2},\quad\text{for all }\lambda>0.$ (1.13)
Moreover, the family of parameters $\{(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})\}_{\lambda>0}$ admits a subsequence that converges to a minimal-norm minimizer of $\mathcal{J}_{0}$.
Remark 1.5.
Several remarks are in order:
  • The proof of Theorem 1.2 can be found in Section 6. It relies on the fact that a minimizer of $\mathcal{J}_{0}$ is known explicitly, which is possible due to the interpolation result from Theorem 1.1. However, Theorem 1.2 is not tight to the 2-width MLP used in Theorem 1.1; it can be applied to any architecture (or discrete dynamical system) whose input-output map is continuous with respect to the parameters.
  • This result shows that, as the regularization parameter $\lambda$ tends to zero, the trained network tends to interpolate the dataset while remaining uniformly bounded in parameter norm. In particular, it guarantees that any accumulation point of the sequence of minimizers corresponds to a minimum-norm interpolant. Indeed, this holds given that the arguments of Theorem 1.2 apply to any minimizer of $\mathcal{J}_{0}$, and in particular to those of minimal norm.This has practical implications in training: it justifies the use of small values of $\lambda$ for training since this pushes the model to fit the training data exactly, with finite values of the parameters, close to the optimal ones.
  • In general, we cannot guarantee that $(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})=(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$. This is a consequence of the lack of convexity of the functionals $\mathcal{J}_{\lambda}$ for $\lambda\geq 0$.
[…]
  • Due to Corollary 1.1, when we consider the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}\subset B_{R_{x}}^{d}(0)\times B_{R_{y}}^{m}(0)$, we can provide an explicit estimate for the optimal value of $\mathcal{J}_{\lambda}$ in terms of $N$, $M$, $R_{x}$, and $R_{y}$. Namely, in the case $m=1$, combining Corollary 1.1 with Theorem 1.2, we obtain
    $\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\leq\lambda C(1+R_{x}^{2}N+R_{x}^{2}N^{2}M+R_{y}^{2}M^{2}).$ (1.14)
    However, since the parameters provided in Theorem 1.1 are constructed to handle worst-case scenarios, the bound in (1.14) may be conservative. In practical situations, significantly tighter estimates may be achieved.
Remark 1.6 (Practical application of Theorem 1.2).
These results, and in particular (1.14), facilitate defining stopping criteria for the training process. Indeed, the optimal value of $\mathcal{J}_{\lambda}$ is always bounded above by $\lambda\left|\!\left|\!\left|(\mathcal{W}^{L}_{*},\mathcal{B}^{L}_{*})\right|\!\right|\!\right|_{2}^{2}$. Therefore, for any fixed $\lambda>0$ – even for small values of $\lambda$ – the optimizer can be designed to terminate once the value of $\mathcal{J}_{\lambda}$ falls close to this explicit threshold.
In practice, neural networks are frequently trained using activation functions that do not satisfy the properties in Remark 1.3. Then, constructing or estimating the parameters assuring classification becomes challenging.
This motivates the following question: Can the knowledge of parameter norms that ensure exact classification for a given activation function, for instance the ReLU, still provide useful information to the regularized functional for a perturbed activation function?
To analyze this point, we consider the following neural network
$\begin{cases}\hat{x}^{j}_{i}=\hat{\bm{\sigma}}_{j}(W_{j}\hat{x}^{j-1}_{i}+b_{j}),\quad&\text{for }j\in\{1,\dots,L\},\\ \hat{x}^{0}_{i}=x_{i},\end{cases}$ (1.15)
where each $\hat{\bm{\sigma}}_{j}:\mathbb{R}^{d_{j}}\rightarrow\mathbb{R}^{d_{j}}$ is a continuous activation function, not necessarily the ReLU. Denote by $\hat{\phi}({\mathcal{W}}^{L},{\mathcal{B}}^{L},\cdot)$ the input-output map associated with (1.15). We then train this new model:
$\min_{({\mathcal{W}}^{L},{\mathcal{B}}^{L})}\left\{\hat{\mathcal{J}}_{\lambda}({\mathcal{W}}^{L},{\mathcal{B}}^{L})=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L},{\mathcal{B}}^{L})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\hat{\phi}({\mathcal{W}}^{L},{\mathcal{B}}^{L},x_{i}),y_{i}\right)\right\},$ (1.16)
where $\operatorname{loss}(\cdot,\cdot)$ is a continuous loss function, as before. Thanks to the continuity of $\operatorname{loss}$ and $\hat{\bm{\sigma}}_{j}$, and the coercivity of $\hat{\mathcal{J}}_{\lambda}$, we can guarantee the existence of a minimizer $(\hat{\mathcal{W}}^{L}_{\lambda},\hat{\mathcal{B}}^{L}_{\lambda})$ for every $\lambda>0$. Moreover, the following result holds.
Theorem 1.3.
Let $d,\,N,\,m\in\mathbb{N}$ with $d,\,N,\,m>1$, and let $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\mathbb{R}^{m}$. Let $(\hat{\mathcal{W}}_{\lambda}^{L},\hat{\mathcal{B}}_{\lambda}^{L})$ be a minimizer of $\hat{\mathcal{J}}_{\lambda}$, and let $(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$ be a minimizer of $\mathcal{J}_{0}$ for the ReLU. Define $R_{0}:=\max_{i\in\{1,\dots,N\}}\|x_{i}\|$, and assume that
$\nu_{j}:=\sup_{z\in B^{d_{j}}_{R_{j}}(0)}\|\hat{\bm{\sigma}}_{j}(z)-\bm{\sigma}_{j}(z)\|<\infty,\quad\text{with}\quad R_{j}:=\|W_{j}\|_{2}R_{j-1}+\|b_{j}\|_{2},\quad\text{for }j\in\{1,\dots,L\},$ (1.17)
$R_{j}$ being defined through the optimal parameters of the ReLU model whose values we estimated. Set $\nu:=(\nu_{1},\dots,\nu_{L})$.
Then, for every $\lambda>0$, we have
$\hat{\mathcal{J}}_{\lambda}(\hat{\mathcal{W}}^{L}_{\lambda},\hat{\mathcal{B}}^{L}_{\lambda})\leq\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2}+\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L}),$
where $\mathscr{A}_{\operatorname{loss}}$ is a nonnegative function depending on the loss function $\operatorname{loss}$, and satisfies $\mathscr{A}_{\operatorname{loss}}(\nu,\cdot,\cdot)\rightarrow 0$ as $\|\nu\|_{2}\rightarrow 0$.
The proof of Theorem 1.3 can be found in Section 6.
Remark 1.7.
Several observations are worth:
  • The proof consists on analyzing the deviation between the input-output maps for both activation functions, using arguments similar to those used in Theorem 1.2. The radii $R_{j}$ introduced in (1.17) are key to estimating the deviation between activations at each layer.
  • Observe that the only assumption imposed on $\hat{\bm{\sigma}}_{j}$ is continuity. No further structural properties are required.
  • We have assumed that both architectures are identical, i.e., data evolve through layers of the same dimensions. This assumption is not strictly necessary. When the dynamics (1.15) evolve along a different architecture, with $\hat{d}_{j}\neq d_{j}$, one can always extend the networks by zero-padding to the maximal dimension $\max\{d_{j},\hat{d}_{j}\}$ and proceed to a stability analysis. A similar estimate then follows, now involving a modified term $\hat{\mathscr{A}}_{\operatorname{loss}}$, depending on the padded architecture.
  • In the particular case $\operatorname{loss}(x,y)=\|x-y\|^{2}_{2}$, we obtain the explicit estimate
    $\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})=\displaystyle\begin{cases}\displaystyle 2\|\nu\|_{2}^{2}\left(\frac{\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2L}-1}{\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2}-1}\right)&\text{if }\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2}\neq 1,\\ 2\|\nu\|_{2}^{2}L&\text{if }\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2}=1.\end{cases}$
The previous result is particularly relevant when, instead of considering a fixed activation function $\hat{\bm{\sigma}}$, for instance the ReLU, we study a family of continuous activation functions $\hat{\bm{\sigma}}_{\varepsilon}$ parametrized by $\varepsilon>0$, for instance a regularization of the ReLU. In this case, we can again formulate a training problem, now depending on the parameter $\varepsilon>0$. Namely, we consider
$\min_{({\mathcal{W}}^{L},{\mathcal{B}}^{L})}\left\{{\mathcal{J}}_{\lambda,\varepsilon}({\mathcal{W}}^{L},{\mathcal{B}}^{L})=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L},{\mathcal{B}}^{L})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\phi_{\varepsilon}({\mathcal{W}}^{L},{\mathcal{B}}^{L},x_{i}),y_{i}\right)\right\},$ (1.18)
where $\phi_{\varepsilon}$ denotes the input-output map of the dynamics (1.15) when $\hat{\bm{\sigma}}$ is replaced by $\hat{\bm{\sigma}}_{\varepsilon}$.
By similar arguments, one can ensure the existence of a minimizer of (1.18) for every $\varepsilon,\,\lambda>0$. We denote by $(\hat{\mathcal{W}}^{L}_{\lambda,\varepsilon},\hat{\mathcal{B}}^{L}_{\lambda,\varepsilon})$ the corresponding minimizer.
A particularly interesting case arises when we define the family $\hat{\sigma}_{\varepsilon}:\mathbb{R}\rightarrow\mathbb{R}$ by
$\hat{\sigma}_{\varepsilon}(x)=\frac{x}{2}\left(1+\operatorname{erf}\left(\frac{x}{\varepsilon\sqrt{2}}\right)\right),$ (1.19)
so that $\hat{\sigma}_{1}(x)=\operatorname{GELU}(x)$, the so-called Gaussian Error Linear Unit (see Figure 6). Let $\hat{\bm{\sigma}}_{\varepsilon}:\mathbb{R}^{d_{j}}\rightarrow\mathbb{R}^{d_{j}}$ denote the vector-valued extension of $\hat{\sigma}_{\varepsilon}$ (as in (1.1)). Then, we obtain the following corollary.
Figure 6. Illustration of the interpolation between ReLU and GELU via $\hat{\sigma}_{\varepsilon}(x)$ for different values of $\varepsilon$. The case $\varepsilon=0$ corresponds to ReLU, and $\varepsilon=1$ to GELU.
Corollary 1.4.
Let $d,\,N,\,m\in\mathbb{N}$ with $d,\,N,\,m>1$, and let $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\mathbb{R}^{m}$. Let $({\mathcal{W}}_{\lambda,\varepsilon}^{L},{\mathcal{B}}_{\lambda,\varepsilon}^{L})$ be a minimizer of ${\mathcal{J}}_{\lambda,\varepsilon}$, and let $(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})$ be a minimizer of $\mathcal{J}_{0}$. Then, under the assumptions of Theorem 1.3, we have
$\limsup_{\varepsilon\rightarrow 0}\mathcal{J}_{\lambda,\varepsilon}({\mathcal{W}}^{L}_{\lambda,\varepsilon},{\mathcal{B}}^{L}_{\lambda,\varepsilon})\leq\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2},$
for every $\lambda>0$.
Remark 1.8.
Corollary 1.4 follows directly from Theorem 1.3, observing that $\sigma_{\varepsilon}\rightarrow\sigma$ in $C(\mathbb{R})$ as $\varepsilon\rightarrow 0$, where $\sigma_{\varepsilon}$ is defined in (1.19), and $\sigma$ denotes the ReLU activation function. The proof is provided in Section 6.
@@ 1.3.2. Implications for Neural Network Training
We now analyze the property of universal approximation.
@@
$\|\phi(\mathcal{W}^{\mathcal{L}},\mathcal{B}^{\mathcal{L}},\cdot)-f(\cdot)\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon.$ (1.20)
@@
$\mathcal{L}\leq C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{d}\varepsilon^{-d},$ (1.21)
@@
Theorem 1.4 ensures the existence of a neural network with a fixed width $d+1$ neurons and sufficient depth approximating any function in $L^{p}(\Omega;\mathbb{R}_{+})$. As illustrated in the strategy of proof below, the neural network is constructed using geometrical arguments. Therefore, Theorem 1.4 not only merely asserts the existence of parameters, but also provides explicit ones.
@@
Step 1. Let $f\in L^{p}(\Omega;\mathbb{R}_{+})$ be a given function and $\varepsilon>0$. Denote by $\mathcal{C}$ the smallest hyperrectangle containing $\Omega$, oriented according to the axes of the canonical basis of $\mathbb{R}^{d}$. We extend $f$ by zero into $\mathcal{C}$. Then, we construct a particular simple function $f_{h}$ that approximates $f$. For its construction, we consider a family $\mathcal{H}_{h}$ of hyperrectangles of size $h>0$ such that $\mathcal{H}_{h}\cup G_{h}^{\delta}=\mathcal{C}$, where $G_{h}^{\delta}$ is a grid with thickness $\delta\leq h^{1+p}$ that satisfies $m_{d}(G_{h}^{\delta})\rightarrow 0$ as $\delta\rightarrow 0$. This is illustrated in Figure 7. Then, $f_{h}$ is defined as the average value of the function $f$ on each hyperrectangle of the family $\mathcal{H}_{h}$. With this simple function, we can guarantee that there exists $h_{1}>0$ such that for every $h<h_{1}$, we have $\|f-f_{h}\|_{L^{p}(\mathcal{C};\mathbb{R}_{+})}<\varepsilon/2$. We denote by $N_{h}$ the number of hyperrectangles in $\mathcal{H}_{h}$.
Figure 7. (A) Level sets of the paraboloid $f(x,y)=x^{2}+y^{2}$ on $\Omega\subset\mathbb{R}^{2}$. The rectangle represents the set $\mathcal{C}$. (B) The function $f$ extended by zero to $\mathcal{C}$. (C) Level sets of $f$ on $\mathcal{H}=\mathcal{C}\setminus G_{\delta}^{h}$, $G_{\delta}^{h}$ being the white mesh.
Step 2. Then, we construct a neural network $\phi^{\mathcal{L}}$ with a width $d+1$ and depth $\mathcal{L}$, ensuring the existence of $h_{2}>0$ such that for all $h<h_{2}$, we have $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(\mathcal{C};\mathbb{R}_{+})}<\varepsilon/2$. This is done in two steps:
Step 2.1. Let $\mathcal{H}^{E}_{h}$ be the subset of hyperrectangles in $\mathcal{H}_{h}$ that are closest to the edges of $\mathcal{C}$, with $N^{E}_{h}$ denoting the number of hyperrectangles in $\mathcal{H}^{E}_{h}$. For each hyperrectangle $H\in\mathcal{H}^{E}_{h}$, we construct a two-layer neural network with an input-output map $\phi_{1}^{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. This map, which has a width $d+1$, drives $H$ to a single point in $\mathbb{R}^{d}$ while mapping the remaining hyperrectangles in $\mathcal{H}_{h}$ to distinct, non-overlapping hyperrectangles in different locations. Here, the chosen parameters ensure the injectivity of the neural network with respect to the hyperrectangles in $\mathcal{H}_{h}$.
@@
Step 2.2 Then, we apply Theorem 1.1 to construct an input-output map $\phi^{L}$ driving the points $\{x_{i}\}_{i=1}^{N_{h}}$ to the $M_{h}$ values of $f_{h}$, ensuring that the composition $\phi^{\mathcal{L}}:=(\phi^{L}\circ\phi^{2N^{E}_{h}})$ equals $f_{h}$ on $\mathcal{H}_{h}$, and therefore $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(\mathcal{H}_{h};\mathbb{R}_{+})}=0$. Additionally, we estimate the error introduced by the neural network on $G_{h}^{\delta}$ and show that, for a fixed $h<1$, the norm $\|\phi^{\mathcal{L}}\|_{L^{\infty}(G_{h}^{\delta};\mathbb{R}_{+})}$ is bounded. This estimation allows us to ensure that there exists a $h_{2}>0$ such that for all $h<h_{2}$, we have $\|f_{h}-\phi^{\mathcal{L}}\|_{L^{p}(G_{h}^{\delta};\mathbb{R}_{+})}<\epsilon/2$.
Step 3. As a consequence of the triangle inequality and by choosing $h<\min\{h_{1},h_{2}\}$, we obtain $\|f-\phi^{\mathcal{L}}\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon$.
Step 4. Due to the explicit construction of the neural network $\phi^{\mathcal{L}}$, the depth $\mathcal{L}$ can be estimated in terms of $h$. Additionally, assuming $f$ to be in $W^{1,p}$, explicit estimates for $h_{1}$ and $h_{2}$ can be obtained in terms of $\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}$, $\varepsilon$, $p$ and $d$ (see [11] and its application to isotropic partitions), and therefore conclude (1.21).
@@
  • Step 2, involving the neural network approximation, is the most challenging one in the proof.
  • As shown in Section 2, each neuron in the neural network represents a hyperplane. In step 2.1, $d+1$ neurons are necessary because this is the number of hyperplanes required to separate a single edge hyperrectangle in $\mathcal{H}^{E}_{h}$. For instance, in the 2-dimensional case, a hyperrectangle on the left edge of $\mathcal{C}$ has only one adjacent hyperrectangle above, one below, and one to its right (there are no hyperrectangles outside $\mathcal{C}$). Thus, only $d+1=3$ hyperplanes are necessary to separate this hyperrectangle from the others, as opposed to the case where a hyperrectangle is in the interior of $\mathcal{C}$, where four hyperrectangles would surround it, and it would be necessary to use $42d=$ 4 hyperplanes. These $d+1$ hyperplanes, in particular, ensure the compression of the separated hyperrectangle into a point.
  • The set $\mathcal{H}^{E}_{h}$ allows us to define a neural network with $2N^{E}_{h}$ layers, mapping the $N_{h}$ hyperrectangles of $\mathcal{H}_{h}$ to $N_{h}$ distinct points. Focusing on $\mathcal{H}^{E}_{h}$ is essential, as it reduces the number of required layers. Mapping each hyperrectangle of $\mathcal{H}_{h}$ individually would result in a neural network with significantly more layers, i.e., $N_{h}\gg N^{E}_{h}$.
  • The neural network in Theorem 1.4, defined as $\phi^{\mathcal{L}}=\phi^{L}\circ\phi^{2N_{h}^{E}}$ requires a width $d+1$. However, only the first $2N^{E}_{h}$ layers require this width. The width of the $L$ remaining layers can be reduced to $2$ (as a consequence of Theorem 1.1).
  • The neural network of width $2$ from Theorem 1.1 maps the compressed hyperrectangles to their corresponding labels and requires $L=2N_{h}+4M_{h}-1$ layers, where $M_{h}$ is the number of distinct values taken by the approximating simple functions $f_{h}$ (which act as labels). In particular, $M_{h}\leq N_{h}$, and $N_{h}$ can be estimated in terms of $h$ and $\delta$. Moreover, since $N^{E}_{h}\leq N_{h}$ and $M_{h}\leq N_{h}$, the depth of the neural network $\phi^{\mathcal{L}}$ is bounded by $\mathcal{L}=2N^{E}_{h}+2N_{h}+4M_{h}-1\leq 8N_{h}-1$. Finally, following Step 4, we can estimate the total depth.
Remark 1.11 (An equivalent statement of Theorem 1.4).
Observe that Theorem 1.4 asserts that for every $\mathcal{L}>0$, there exist parameters $\mathcal{W}^{\mathcal{L}}$ and $\mathcal{B}^{\mathcal{L}}$ such that the input-output map of (1.2) with $w_{\max}=d+1$ satisfies
$\|\phi^{\mathcal{L}}(\mathcal{W}^{\mathcal{L}},\mathcal{B}^{\mathcal{L}},\cdot)-f(\cdot)\|_{L^{p}(\Omega;\mathbb{R}_{+})}\leq C^{\prime}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}\mathcal{L}^{-1/d},$
where $C^{\prime}$ is a positive constant depending on $m_{d}(\Omega)$, $d$, and $p$.
It is important to mention that in the absence of a finite $W^{1,p}$-norm, a function $f\in L^{p}(\Omega;\mathbb{R}_{+})$ may exhibit arbitrarily rapid oscillations, making any nontrivial approximation rate unattainable. Therefore, the finiteness of the $W^{1,p}$-norm is essential.